OpenStudy (anonymous):

please help . . . A painter earns $15 per hour. What is the minimum number of hours he must work to earn at least $200? Write an inequality to represent this situation and solve. A) 15h *greater than or equal to* 200; h *greater than or equal to* 13 and one third B) 15h *less than or equal to* 200; h *less than or equal to* 13 and one third C) 200h *greater than or equal to* 15; h *greater than or equal to* three fortieths D) 200h *less than or equal to* 15; h *less than or equal to* 3 fortieths

OpenStudy (anonymous):

Well you could subtract 15 dollars a bunch of times from 200 dollars to figure out the number of hours it takes to make 200 bucks. 200-15= 185 one hour 185-15= 170 two hours 170-15= 155 three hours You could keep going until you can't subtract any more 15 dollar amounts or you could recognize that repeated subtraction is just division. So 200 divided by 15 is the number of hours it takes to make exactly 200 dollars. So if the painter wants to make more than 200 dollars, the painter has to work more hours than 200 divided by 15 figure.